OpenStudy (anonymous):

Let the function h be defined by h(x)=14+9m, x sq./4. If h(2m)=9m, what is one possible value of m? If h(2m)=

OpenStudy (chihiroasleaf):

what do you mean by x sq./4 ?

OpenStudy (anonymous):

x to the power of 2 divided by 4

OpenStudy (chihiroasleaf):

hhmm..., it's confusing... :D so, define a function h with \[h(x) = 14 + 9m\] if \[h(2m) = 9m\] what is m? right? so what is \[\frac{ x ^{2} }{ 4 } ?\]

OpenStudy (anonymous):

wouldn't you just plug in 2 from 2m?

OpenStudy (chihiroasleaf):

are you sure that you don't mistype the question?

OpenStudy (anonymous):

Let the function h be defined by h(x)=14+x to the power of 2 divided by 4. If h(2m)=9m, what is one possible value of m?

OpenStudy (anonymous):

the equation is 14+ 2m squared/4

OpenStudy (chihiroasleaf):

ahh..., so, \[h(x) = 14 + \frac{ x ^{2} }{ 4 }\] if \[h(2m) = 9m\] plug in x = 2m into the definition of the function

OpenStudy (anonymous):

7 +m squared=h(x)

OpenStudy (chihiroasleaf):

be careful, it's x = 2m, so it's (2m)^2

OpenStudy (anonymous):

14+2m squared divided by 4

OpenStudy (anonymous):

IF I cross cancel, I get 7+2m squared divided by 2

OpenStudy (anonymous):

that gives me 7+ m squared=h(x)

OpenStudy (chihiroasleaf):

how do you get 7 + 2m squared ?

OpenStudy (anonymous):

14+2x squared divided b6 4

OpenStudy (anonymous):

I cross cancel with 14 and 4, which gives me 7 and 2

OpenStudy (chihiroasleaf):

it's plus, not multiply, right? you cannot use 'cross cancel'

OpenStudy (chihiroasleaf):

*addition I mean

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

then 14+ m squared/2

OpenStudy (anonymous):

=h(x)

OpenStudy (chihiroasleaf):

be careful, it's (2m)^2 = 2^2 x m^2 = 4m^2

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

14+ m sq.=h(x)?

OpenStudy (chihiroasleaf):

yes, it uses the properties of exponent (ab)^m = a^m x b^m 14 + m^2 = h(2m) = 9m m^2 - 9m + 14 = 0 can you solve this quadratic equation?

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

m=2 or 7

OpenStudy (chihiroasleaf):

yup :)

OpenStudy (anonymous):

next problem?